OpenStudy (pottersheep):

I don't understand the "u" method properly? Pleaseee help. x^1/2 - 3x^1/4 + 2 = 0 What would "u" be? x^1/2 or x^1/4 ?

OpenStudy (pottersheep):

Apparently its supposed to be u = x^1/4 , why would u^2 = 1/2?

OpenStudy (anonymous):

what are you trying to do? solve for the roots?

OpenStudy (pottersheep):

Nvm, here's a medal anyways :)